# A New Formulation for Zeroth-Order Optimization of Adversarial EXEmples in Malware Detection

### v1

@@ Abstract
Machine learning malware detectors are vulnerable to adversarial EXEmples, i.e. carefully-crafted Windows programs tailored to evade detection. Unlike other adversarial problems, attacks in this context must be functionality-preserving, a constraint which is challenging to address. As a consequence heuristic algorithms are typically used, that inject new content, either randomly-picked or harvested from legitimate programs. In this paper, we show how learning malware detectors can be cast within a zeroth-order optimization framework which allows to incorporate functionality-preserving manipulations. This permits the deployment of sound and efficient gradient-free optimization algorithms, which come with theoretical guarantees and allow for minimal hyper-parameters tuning. As a by-product, we propose and study ZEXE, a novel zero-order attack against Windows malware detection. Compared to state-of-the-art techniques, ZEXE provides drastic improvement in the evasion rate, while reducing to less than one third the size of the injected content.
@@ 1 Introduction
Every week,  7M malicious software, namely known as malware, are blocked by detection engines on VirusTotal. [^1] https://www.virustotal.com/gui/stats To cope with this sheer amount of threats, we assisted to the rise of machine-learning Windows malware detectors with superhuman performances, becoming the de-facto standard approach in this domain [1, 29, 35, 17]. However, recent work proved the existence of adversarial EXEmples, carefully-manipulated Windows malware that bypass detection [13]. Differently from adversarial examples, which are computed by summing well-crafted noise to pixels of input images, these EXEmples must adhere to strict domain constraints imposed by the Windows PE file format, [^2] https://learn.microsoft.com/en-gb/windows/win32/debug/pe-format which precisely specifies how programs are stored as files. Hence, attackers must abandon additive manipulations in favor of practical manipulations that perturbs the structure of programs without causing corruptions while maintaining the original functionality [12]. As a result, the literature thrives with adversarial EXEmples attacks [13, 24, 34, 2, 21], with a strong focus on gradient-free techniques that inject new content, either randomly-picked or harvested from legitimate programs known as benignware, to boost the likelihood of evasion [12].
While shown to be effective, the inclusion of practical manipulations increase the complexity of formally defining the optimization problem. Thus, none of the proposed gradient-free attacks provide theoretical guarantees since they are not supported by a formulation that permits such proofs. Also, albeit being the most powerful strategies also in black-box scenarios where the target can only be reached through queries, current content-injection attacks [12, 2, 34] rely on unbounded perturbations, causing an uncontrolled growth in size of manipulated programs without the possibility of fixing it a priori. Lastly, the choice of the injected content deeply influences the evasiveness, since it presents itself as an hyper-parameter of the optimization algorithm, thus requiring proper tuning.
@@
Hence, in this paper we overcome these limitations by proposing a novel formulation of adversarial EXEmples that defines practical manipulations as: (i) an initialization step that tackles the generation of fixed-size perturbation inside programs, and (ii) the rewriting step that overwrites the bytes selected by the initialization (Sec. 3.1). With this trick, we can write a non-convex, non-smooth optimization problem directly in $\mathbb{R}^{d}$, while preserving the functionality in the original input space, being also solvable with algorithms backed with provable guarantees. We then propose ZEXE, a zeroth-order optimization algorithm based on prior work [31] that leverages finite difference to estimate gradients and provides theoretical guarantees for non-convex non-smooth functions (Sec. 3.2). Moreover, we propose an adaptive strategy to select the parameters avoiding manual tuning. We test ZEXE against GAMMA [12] one of the most-powerful content-injection attack, showing that, with the same amounts of queries, our technique requires half of the perturbation size to drastically increase the evasion rate against two state-of-the-art machine learning Windows malware detectors (Sec. 4).
@@ 2 Background
Adversarial EXEmples. Machine learning models are vulnerable to adversarial examples, carefully-crafted test time samples that coerce the output of classification towards answers decided by attackers [5]. While these threats have been widely formulated on the image domain, recent work also showed their efficacy in more complex scenarios like malware detection, where attackers create adversarial EXEmples, Windows programs tailored to evade detection [12, 13]. These are computed as the solution of a minimization problem:
@@
where $L$ is a loss function of choice that quantifies the error committed by a malware detector $f$ in assigning the class of legitimate programs $y$. This error is computed on a malware sample $z\in Z$ (where $Z\in[0,255]^{*}$ is the space of all programs stored as strings of bytes) perturbed with the practical manipulation function $\rho:Z\times\Delta\rightarrow Z$ which is functionality-preserving by design. This means that input programs perturbed with $\rho$ are still functional without alteration in their execution, without requiring the inclusion of hard-constraints in Eq. 1. We remark that these constraints are not investigated in the usual domain of images, since the careless perturbation of bytes inside programs cause corruptions, rendering them unsable. Lastly, the $\rho$ function injects a manipulation $\delta\in\Delta$, being $\Delta$ the space of the perturbation, which can be differently formalized [13].
Structured Finite-Difference Method. To compute solutions of optimization problems without exploiting gradients, it is possible to rely on zeroth-order methods that leverage finite-difference algorithms [20, 26], which substitute the gradient of an objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ with a surrogate $g$ built through finite differences on a set of random directions $(p^{(i)})_{i=1}^{\ell}$ with $\ell>0$. These directions can be either (i) sampled i.i.d. from a distribution (unstructured methods), or (ii) chosen through structural constraints like orthogonality (structured methods). In particular, [3] observed that the latter provide a better approximation of the gradient with respect to unstructured ones. These rely on random orthogonal directions that can be represented as rotations of the first $\ell\leq d$ canonical bases $(e_{i})_{i=1}^{\ell}$. Formally, we set $p^{(i)}=Ge_{i}$ where $G$ belongs to the orthogonal group $O(d)$ defined as
$O(d):=\{G\in\mathbb{R}^{d\times d}\,|\,\det G\neq 0\quad\wedge\quad G^{-1}=G^{\intercal}\}.$
Some methods to generate orthogonal matrices are described in [31, Appendix D]. Let $G\in O(d)$, $h>0$ and $\ell\leq d$, we build a gradient surrogate as
@@
where $\gamma_{k}>0$ is a stepsize. This algorithm has been analyzed in different settings [31]. In particular, to perform the analysis in nonsmooth setting, a smoothing technique is exploited. Let $h>0$, we define a smooth approximation of the objective function
[…]
where $\text{vol}(\mathbb{B}^{d})$ denotes the volume of the unitary $\ell_{2}$ ball $\mathbb{B}^{d}:=\{x\in\mathbb{R}^{d}\,|\,\|x\|_{2}\leq 1\}$. In particular, [4] showed that $f_{h}$ is differentiable even when $f$ is not. By [31, Lemma 1], we have that $\mathbb{E}_{G}[g_{(G,h)}(x)]=\nabla f_{h}(x)$. This allows to interpret the iteration (3) as Stochastic Gradient Descent on the smooth approximation of the objective function $F_{h_{k}}$.
@@ 3 Zeroth-Order Optimization of Adversarial EXEmples
In this section we propose our novel formulation for adversarial EXEmples (Sec. 3.1), followed by our new zeroth-order optimization algorithm ZEXE (Sec. 3.2) and its theoretical guarantees (Sec. 3.3).
@@ 3.1 Formulation of Adversarial EXEmples attacks
We propose a novel formulation for Adversarial EXEmples crafting problem that permits the application of well-established zeroth-order optimization algorithm. The idea consists in modifying the definition of the manipulation $\delta$ in Eq. 1. First, given an initial manipulation $\delta\in\Delta$, an target program $z\in Z$ and a manipulation mapping $\rho:Z\times\Delta\to Z$, we can construct a manipulated program as $\bar{z}=\rho(z,\delta)$. This initial manipulation can be interpreted as the initialization of our attack. Note that the introduction of an initial manipulation allows us to view the program as composed by two parts: the effective program $z$ which cannot be modified and the manipulation $\delta$ which can be perturbed. To effectively adjust the manipulation, we define an encoding function $\psi:\Delta\to\mathbb{R}^{d}$ that maps byte values into the corresponding real numbers. Once the manipulation is encoded using this function, we can apply our perturbation just summing it to the encoded manipulation. Subsequently, to revert the perturbed data back to its original byte format, we define a decoder function $\xi:\mathbb{R}^{d}\to\Delta$. This decoder function serves as the counterpart to the encoder, effectively mapping the perturbed real vector back to its byte representation. Finally, we apply the manipulation mapping obtaining a perturbed program. We, thus, formulate the problem of creating the perturbation to be applied to a program $z$ to compute an adversarial EXEmples as:
@@
Note that with this formulation we can use any zeroth-order optimization algorithm working on $\mathbb{R}^{d}$.
@@ 3.2 ZEXE: Adaptive zeroth-Order attack against Windows Malware Detectors
In this section, we propose a zeroth-order algorithm to solve the problem formulated in Eq. 5. Our approach extends the random search algorithm by merging random sampling with the finite-difference method. Given a sampling region $V\subseteq[0,255]^{d}$, the algorithm samples a random initial guess (initialization), it explores ’near’ configurations (local exploration) and refines such a guess with a finite-difference strategy (refinement).
Initialization. Given a program $z$, the algorithm begins by evaluating the objective function $F_{z}:R^{d}\to R$ at the initial manipulation $v_{0}\in V$. The function value and the manipulation are stored as the current best evaluation and iteration found respectively i.e.,
@@
Local Exploration. Then, at each iteration $k\in\mathbb{N}$, the algorithm samples a random orthogonal matrix $G_{k}$ from the orthogonal group $O(d)$ and it builds the following sets
$V_{k}:=\{(v_{k}+h_{k}G_{k}e_{i})\}_{i=1}^{\ell}\cup\{v_{k},v_{k}^{\text{best}}\}\quad\text{and}\quad D_{k}:=\{F_{z}(v_{k}+h_{k}G_{k}e_{i})\}_{i=1}^{\ell}\cup\{F_{z}(v_{k}),F_{k}^{\text{best}}\}.$
Note that this operation requires only $\ell+1$ function evaluations since $F_{k}^{\text{best}}=F_{z}(v_{k}^{\text{best}})$. Then, we update the current best iterate and function value comparing the current best value with the function values computed. Formally,
@@
Note that our surrogate is different from Eq. 2. Indeed, our approximation can be computed in $\ell+1$ function evaluations instead of $2\ell$. We decided to use this because we observed that it permits to get better performance in practice. The manipulation is then updated according to the following iteration
$v_{k+1}=\left\{\begin{array}[]{ll}v_{k}-\gamma_{k}g_{(G_{k},h_{k})}(v_{k})&\text{if }|g_{(G_{k},h_{k})}(v_{k})|^{2}>\omega,\\ \text{sample }v\in V\subseteq[0,255]^{d}\text{ at random}&\text{otherwise}.\end{array}\right.$ (8)
Note that for $\omega=0$, Algorithm 1 recovers the standard finite difference iteration analyzed in different works with different gradient approximations  [26, 16, 15]. Choosing $\omega>0$ allows for exploration behavior. Intuitively, if $g$ is a good approximation of the gradient of $F_{z}$, a small norm indicates that the algorithm is approaching a stationary point. To avoid getting stuck, we resample a new point allowing to explore a different region of the space. The algorithm can be summarized with the following pseudo-code.
Algorithm 1 ZEXE: zeroth-order optimization algorithm for crafting adversarial EXEmples
@@
$v_{k+1}=\left\{\begin{array}[]{ll}v_{k}-\gamma_{k}g_{k}&\text{if }\|g_{k}\|^{2}\geq\omega\\ \text{sample }v\in V\text{ uniformely at random}&\text{otherwise}\end{array}\right.$
@@
Choice of the parameters. Algorithm 1 depends on different parameters namely the stepsizes $\gamma_{k}$, the smoothing parameters $h_{k}$, the number of directions $\ell$ and $\omega$. Since parameter tuning can be hard to perform in practice, we propose the following adaptive stepsize and smoothing parameters. Given $0<\underline{\gamma}\leq\bar{\gamma}$, we define
$\gamma_{k}=\max(\underline{\gamma},\min(F_{z}(v_{k})/\|g_{(G_{k},h_{k})}(v_{k})\|^{2},\bar{\gamma})).$ (9)
Moreover, given $\underline{h}>0$ and $h_{0}\geq\underline{h}$, we define the following sequence of smoothing parameter

### v2

@@ Abstract
Machine learning malware detectors are vulnerable to adversarial EXEmples, i.e., carefully-crafted Windows programs tailored to evade detection. Unlike other adversarial problems, attacks in this context must be functionality-preserving, a constraint that is challenging to address. As a consequence, heuristic algorithms are typically used, which inject new content, either randomly-picked or harvested from legitimate programs. In this paper, we show how learning malware detectors can be cast within a zeroth-order optimization framework, which allows incorporating functionality-preserving manipulations. This permits the deployment of sound and efficient gradient-free optimization algorithms, which come with theoretical guarantees and allow for minimal hyper-parameters tuning. As a by-product, we propose and study ZEXE, a novel zeroth-order attack against Windows malware detection. Compared to state-of-the-art techniques, ZEXE provides improvement in the evasion rate, reducing to less than one third the size of the injected content.
footnotetext: This is the preprint version of the following IEEE-published article: M. Rando, L. Demetrio, L. Rosasco and F. Roli, "A New Formulation for Zeroth-Order Optimization of Adversarial EXEmples in Malware Detection," in IEEE Transactions on Information Forensics and Security, vol. 21, pp. 506-515, 2026, doi: 10.1109/TIFS.2025.3648867.
@@ 1 Introduction
Every week, 7M malicious software, known as malware, are blocked by detection engines on VirusTotal. [^1] https://www.virustotal.com/gui/stats While most of these are known samples that can be detected through pattern-matching with signatures, we assisted in the rise of machine-learning Windows malware detectors to deal with unknown variants that slip through signature-based detection [2, 31, 38, 18]. However, recent works proved the existence of adversarial EXEmples, carefully-manipulated Windows malware that bypass detection [14]. Differently from adversarial examples, which are computed by summing well-crafted noise to pixels of input images, these EXEmples must adhere to strict domain constraints imposed by the Windows PE file format, [^2] https://learn.microsoft.com/en-gb/windows/win32/debug/pe-format, which precisely specifies how programs are stored as files. Hence, attackers must abandon additive manipulations in favor of practical manipulations that perturb the structure of programs without causing corruption while maintaining the original functionality [12]. As a result, the literature thrives with adversarial EXEmples attacks [14, 27, 37, 1, 22], with a strong focus on gradient-free techniques that inject new content, either randomly-picked or harvested from legitimate programs known as benignware, to boost the likelihood of evasion [12]. While shown to be effective, the inclusion of practical manipulations increases the complexity of formally defining the optimization problem. Thus, none of the proposed gradient-free attacks provides theoretical guarantees since they are not supported by a formulation that permits such proofs. Also, although being the most powerful strategies also in black-box scenarios where the target can only be reached via queries, current content-injection attacks [12, 1, 37] rely on unbounded perturbations, causing an uncontrolled growth in size of manipulated programs without the possibility of fixing it a priori. Lastly, the choice of the injected content deeply influences the evasiveness, since it presents itself as a hyper-parameter of the optimization algorithm, thus requiring proper tuning. Hence, in this paper, we overcome these limitations by proposing a novel formulation of adversarial EXEmples that defines practical manipulations as: (i) an initialization step that tackles the generation of fixed-size perturbations inside programs, and (ii) a rewriting step that overwrites the bytes selected by the initialization (Sec. 3.1). With this trick, we can write a non-convex, non-smooth optimization problem directly in $\mathbb{R}^{d}$, while preserving the functionality in the original input space, being also solvable with algorithms backed by provable guarantees. We then propose ZEXE, a zeroth-order optimization algorithm based on prior work [32] that leverages finite differences to estimate gradients and provides theoretical guarantees for non-convex, non-smooth functions (Sec. 3.2) along with an adaptive strategy to select the parameters, avoiding manual tuning. We test ZEXE against GAMMA [12], one of the most powerful content-injection attacks, showing that, with the same number of queries, our technique requires half of the perturbation size to increase the evasion rate against two state-of-the-art machine learning Windows malware detectors (Sec. 4).
@@ 2 Background
Adversarial EXEmples. Machine learning models are vulnerable to adversarial examples, carefully-crafted test-time samples that coerce the output of classification towards answers decided by attackers [6]. Widely formulated on the image domain, recent work has also shown their efficacy in more complex scenarios like malware detection, where attackers create adversarial EXEmples, Windows programs tailored to evade detection [12, 14] computed as the solution of a minimization problem:
@@
where $L$ is a loss function of choice that quantifies the error committed by a malware detector $f$ in assigning the class of legitimate programs $y$. This error is computed on a malware sample $z\in Z$ (where $Z\in[0,255]^{*}$ is the space of all programs stored as strings of bytes) perturbed with the practical manipulation function $\rho:Z\times\Delta\rightarrow Z$, which is functionality-preserving by design. This means that input programs perturbed with $\rho$ are still functional without alteration of their execution, without requiring the inclusion of hard constraints in Eq. 1. We remark that these constraints are not investigated in the usual domain of images, since the careless perturbation of bytes inside programs causes corruption, rendering them unstable. Lastly, the $\rho$ function injects a manipulation $\delta\in\Delta$, where $\Delta$ is the space of the perturbation, which can be differently formalized [14].
Structured Finite-Difference Method. To compute solutions of optimization problems without exploiting gradients, it is possible to rely on zeroth-order methods that leverage finite-difference algorithms [21, 28], which substitute the gradient of an objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ with a surrogate $g$ built through finite differences over a set of random directions $(p^{(i)}){i=1}^{\ell}$ with $\ell>0$. These directions can be either (i) sampled i.i.d. from a distribution (unstructured methods), or (ii) chosen through structural constraints like orthogonality (structured methods). In particular, [3] observed that the latter provides a better approximation of the gradient compared to unstructured ones. These rely on random orthogonal directions that can be represented as rotations of the first $\ell\leq d$ canonical basis vectors $(e_{i}){i=1}^{\ell}$. Formally, we set $p^{(i)}=Ge_{i}$ where $G$ belongs to the orthogonal group $O(d)$ defined as
$O(d):={G\in\mathbb{R}^{d\times d},|,\det G\neq 0\quad\wedge\quad G^{-1}=G^{\intercal}}.$
Some methods to generate orthogonal matrices are described in [32, Appendix D]. Let $G\in O(d)$, $h>0$ and $\ell\leq d$, we build a gradient surrogate as
@@
where $\gamma_{k}>0$ is a stepsize. This algorithm has been analyzed in different settings [32]. In particular, to perform the analysis in the non-smooth setting, a smoothing technique is exploited. Let $h>0$, we define a smooth approximation of the objective function
$f_{h}(x):=\frac{1}{\text{vol}(\mathbb{B}^{d})}\int f(x+hu)du,$ (4)
where $\text{vol}(\mathbb{B}^{d})$ denotes the volume of the unitary $\ell_{2}$ ball $\mathbb{B}^{d}:={x\in\mathbb{R}^{d},|,|x|2\leq 1}$. In particular, [5] showed that $f_{h}$ is differentiable even when $f$ is not. By [32, Lemma 1], we have that $\mathbb{E}G[g{(G,h)}(x)]=\nabla f_{h}(x)$. This allows us to interpret the iteration (3) as Stochastic Gradient Descent on the smooth approximation of the objective function $F{h_{k}}$.
@@ 3 Zeroth-Order Optimization of Adversarial EXEmples
In this section, we propose our novel formulation for adversarial EXEmples (Sec. 3.1), followed by our new zeroth-order optimization algorithm ZEXE (Sec. 3.2) and its theoretical guarantees (Sec. 3.4).
@@ 3.1 Formulation of Adversarial EXEmples attacks
We propose a novel formulation for Adversarial EXEmples crafting problem that permits the application of well-established zeroth-order optimization algorithm. Such a definition addresses all those scenarios where the model to evaluate is either non-differentiable, or only accessible through scores only. The idea consists of modifying the definition of the manipulation $\delta$ in Eq. 1. First, given an initial manipulation $\delta\in\Delta$, an target program $z\in Z$ and a manipulation mapping $\rho:Z\times\Delta\to Z$, we can construct a manipulated program as $\bar{z}=\rho(z,\delta)$. This initial manipulation can be interpreted as the initialization of our attack. Note that the introduction of an initial manipulation allows us to view the program as composed of two parts: the effective program $z$, which cannot be modified, and the manipulation $\delta$, which can be perturbed. To effectively adjust the manipulation, we define an encoding function $\psi:\Delta\to\mathbb{R}^{d}$ that maps byte values into the corresponding real numbers. Once the manipulation is encoded using this function, we can apply our perturbation just summing it to the encoded manipulation. Subsequently, to revert the perturbed data back to its original byte format, we define a decoder function $\xi:\mathbb{R}^{d}\to\Delta$. This decoder function serves as the counterpart to the encoder, effectively mapping the perturbed real vector back to its byte representation. Finally, we apply the manipulation mapping to obtain a perturbed program. We, thus, formulate the problem of creating the perturbation to be applied to a program $z$ to compute an adversarial EXEmples as:
@@
Note that with this formulation, we can use any zeroth-order optimization algorithm working on $\mathbb{R}^{d}$.
@@ 3.2 ZEXE: Adaptive Zeroth-Order attack against Windows Malware Detectors
In this section, we propose a zeroth-order algorithm to solve the problem formulated in Eq. 5. Our approach extends the random search algorithm by merging random sampling with the finite-difference method. Given a sampling region $V\subseteq[0,255]^{d}$, the algorithm samples a random initial guess (initialization), it explores ’near’ configurations (local exploration), and refines such a guess with a finite-difference strategy (refinement).
Initialization. Given a program $z$, the algorithm begins by evaluating the objective function $F_{z}:R^{d}\to R$ at the initial manipulation $v_{0}\in V$. The function value and the manipulation are stored as the current best evaluation and iteration found, respectively, i.e.,
@@
Local Exploration. Then, at each iteration $k\in\mathbb{N}$, the algorithm samples a random orthogonal matrix $G_{k}$ and builds the following sets
$V_{k}=\{(v_{k}+h_{k}G_{k}e_{i})\}_{i=1}^{\ell}\cup\{v_{k},v_{k}^{\text{best}}\}\quad\text{and}$
$D_{k}=\{F_{z}(v_{k}+h_{k}G_{k}e_{i})\}_{i=1}^{\ell}\cup\{F_{z}(v_{k}),F_{k}^{\text{best}}\}.$
Note that this operation requires only $\ell+1$ function evaluations since $F_{k}^{\text{best}}=F_{z}(v_{k}^{\text{best}})$. Then, we update the current best iterate and function value by comparing the current best value with the function values computed. Formally,
@@
Note that our surrogate is different from Eq. 2. Indeed, our approximation can be computed in $\ell+1$ function evaluations instead of $2\ell$. We decided to use this because we observed that it permits getting better performance in practice. The manipulation is then updated according to the following iteration.
$v_{k+1}=\left\{\begin{array}[]{ll}v_{k}-\gamma_{k}g_{(G_{k},h_{k})}(v_{k})\qquad\text{if }\|g_{(G_{k},h_{k})}(v_{k})\|^{2}>\omega,\\ \text{sample }v\in V\subseteq[0,255]^{d}\text{ at random}\end{array}\right$. (8)
Note that for $\omega=0$, Algorithm 1 recovers the standard finite difference iteration analyzed in different works with different gradient approximations  [28, 17, 16]. Choosing $\omega>0$ allows for exploration behavior. Intuitively, if $g$ is a good approximation of the gradient of $F_{z}$, a small norm indicates that the algorithm is approaching a stationary point. To avoid getting stuck, we resample a new point, allowing us to explore a different region of the space. The algorithm can be summarized with the following pseudo-code.
Algorithm 1 ZEXE: Zeroth-order optimization attack
@@
$v_{k+1}=\left\{\begin{array}[]{ll}v_{k}-\gamma_{k}g_{k}&\text{if }\|g_{k}\|^{2}\geq\omega\\ \text{sample }v\in V\text{ uniformly}&\text{else}\end{array}\right$.
@@
3.3 High-dimensional Perturbations and Exploration Strategies.
The refinement step of Algorithm 1 alternates between moving in the opposite direction of the gradient approximation (exploitation) and performing a random search over a subset of bytes (exploration). As indicated in Eq. 8, the decision of which step to perform depends on the squared norm of the gradient approximation. Specifically, at each iteration $k\in\mathbb{N}$, if $\|g_{k}\|^{2}$ is sufficiently small (i.e., $\|g_{k}\|^{2}<\omega$), the algorithm performs an exploration step. However, when the perturbation size is very large, this approach may become less effective, since a substantial number of function evaluations may be required to identify a suitable perturbation. In such cases, alternative strategies should be considered. To tackle such a setting, we propose two alternatives: ZEXE-Langevin and ZEXE-GA.
ZEXE-Langevin. A possible alternative to random search is to merge exploration and exploitation by adding Gaussian noise to the iterate. For every $k\in\mathbb{N}$, let $\beta_{k}\geq 0$. We replace the iterate of ZEXE (i.e., step $7$ of Algorithm 1) with
$v_{k+1}=v_{k}-\gamma_{k}g_{k}+\sqrt{2\beta_{k}}\xi_{k},\qquad\xi_{k}\sim\mathcal{N}(0,I).$ (9)
The intuition behind this update is that the gradient term encourages exploitation by moving the iterate along the approximated gradient direction, while the Gaussian perturbation introduces an exploration component that allows the algorithm to escape stationary points. Notice that for $\beta_{k}=0$, Eq. 9 reduces to the finite-difference iteration as a special case, while for large $\beta_{k}$ the method approximates random search, since the noise term $\sqrt{2\beta_{k}}\xi_{k}$ dominates the update. Moreover, under suitable assumptions on the target function, this iterate (with a different gradient approximation) has been analyzed in [34], providing theoretical guarantees on its convergence properties.
ZEXE-GA. Another alternative is to employ a Genetic Algorithm (GA) [24] to perform exploration steps. The algorithm maintains a population of $p\in\mathbb{N}_{+}$ perturbations that are randomly initialized. Whenever $\|g_{k}\|^{2}<\omega$, the elements of this population are sequentially used as candidate solutions and evaluated. For each candidate, the corresponding gradient approximation $g_{k}$ is computed. If the condition $\|g_{k}\|^{2}<\omega$ still holds, the algorithm proceeds to the next candidate. Once all candidates in the population have been used, the population is updated through the standard GA operations of selection, mutation, and crossover. When a candidate satisfying $\|g_{k}\|^{2}\geq\omega$ is found, it is used as the current iterate and updated according to Eq. 8.